Circular No. 6259

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CASSIOPEIAE 1995 U. Munari, Asiago Observatory, Padua; and A. Lepardo, R. Pas-

suello, and G. Sostero, Associazione Friulana di Astronomia e Meteorologia (AFAM), communicate: “Photoelectric photometry secured with the 0.45m AFAM reflector on Nov. 4.9 UT yields V = 8.45, B−V = +0.51, U −B = −0.15 (±0.02 mag in all colors). Simultaneous spectra (range 350–900 nm, resolution 0.04 nm) obtained with the Asiago 1.82-m telescope (+ échelle spectrograph + CCD) indicate the object to be a symbiotic nova (similar to PU Vul) and not a classical nova. All absorption lines of the underlying F continuum are narrow. Hα, Hβ, Hγ, the Ca II triplet at 850–866 nm, and many Fe II lines appear in moderate emission, while higher Balmer lines are present in absorption. The profiles consist of sharp absorptions (FWHM 55 km/s) superimposed onto the blue wing of narrow emission (FWHM 95 km/s) components, shifted by ∼ −85 km/s from the latter.”

α CENTAURI B E. Guinan, Villanova University; and S. Messina, University of Cata-

nia, report on the determination of the rotation period of the K1 V component of the α Cen triple system: “The star was observed about twice a week over a three-month interval (May–Aug.) with the IUE satellite (+ LWPHIGH and SWPLOW cameras). These observations show a 20- to 30percent modulation in the chromospheric and transition-region ultraviolet line emissions with a period of 36.8 ± 0.4 days. This period is assumed to be the rotation period, and the modulations of the emission-line strengths arise as large active sites on the star rotate in and out of view. A definite determination of the rotation period of α Cen B is important for testing stellar dynamo theories because the age and physical properties of the α Cen system are known with high precision. Determinations of the rotational periods of α Cen A (G2 V) and α Cen C (= Proxima Cen; dM5e) are also being determined in a similar manner. It is estimated from the amplitudes of the chromospheric lines’ emissions, that α Cen B could show 1- to 2-percent light variations from the rotational modulation by starspots that commonly accompany plage emissions.”

COMET 73P/SCHWASSMANN-WACHMANN 3 Total visual magnitude estimates: Oct. 22.10 UT, 6.3 (J. V. Scotti,

Kitt Peak, AZ, 10×50 binoculars); 28.06, 5.6 (R. Keen, Mt. Thorodin, CO, 7×50 binoculars); Nov. 8.08, 8.1 (A. Hale, Cloudcroft, NM, 0.2-m reflector).

1995 November 11 Daniel W. E. Green